OpenStudy (smurfy14):

81^(3/4)...i know the answer is 27 but how do you get there? do u just divide 81 by 3?

OpenStudy (amistre64):

\[\sqrt[4]{81^3}\] is the only way I can see it

OpenStudy (smurfy14):

whts that mean?

OpenStudy (amistre64):

81^(3/4) = y 3/4 ln(81) = y .... cant see it that way it means you cube 81 and see what it 4roots into

OpenStudy (smurfy14):

ok well how do you solve??

OpenStudy (amistre64):

there is prolly an elegant algorithm for it; but i just have to brute math it: 4rt(81^3) 4rt(531441) = n 531441 = n^4

OpenStudy (amistre64):

then i trial and error it till I reach 27

OpenStudy (anonymous):

81^3/4 = 3^(4*3/4) = 3^3 = 27

OpenStudy (anonymous):

notice that 81 = 9*9 and 9 = 3*3 so 81 = 3*3*3*3 = 3^4

OpenStudy (smurfy14):

how did you get the (4*3/4) part?

OpenStudy (amistre64):

Y=f'(x)(x-xo)+f(xo) perhaps?

OpenStudy (anonymous):

\[81^{3/4} = (3^{4})^{3/4} = 3^{4*3/4} = 3^{3} =27\]

OpenStudy (smurfy14):

amistre your making it too complicating for me lol

OpenStudy (amistre64):

:) srry

OpenStudy (smurfy14):

oohhh ok thank you soooo much dhatraditya :D

OpenStudy (anonymous):

you are welcome. :)

OpenStudy (mathteacher1729):

\[81= 3*3*3*3 \] So we're looking for \[ ((81)^{1/4})^3\] Notice that \[ ((3)^4)^{1/4} = 3\] and \[3^3 = 27\]